OpenStudy (mendicant_bias):

(PDE)(Sturm-Liouville) I don't understand how in this document, in example 6.10, they got the general solution for lambda less than zero. Posted below shortly.

OpenStudy (mendicant_bias):

(Page 3, Example 6.10) http://www.macs.hw.ac.uk/~robertw/F13YT2/chap6.pdf

OpenStudy (mendicant_bias):

How did they get that the general solution for \[y''=k^2y\]equals\[y(x)=A \cosh(kx)+B \sinh(kx)?\]

OpenStudy (mendicant_bias):

(When I say Page 3, I mean 3 in the PDF, not by the page numbers in the document)

OpenStudy (perl):

thats a fun exercise

OpenStudy (perl):

we want to solve the equation y ' ' - k^2 y = 0 you can use power series method

OpenStudy (mendicant_bias):

Is this Froebenius(sp?) Method?

OpenStudy (perl):

i dont think we would use that here. we can use power series. i just did a problem like this yesterday

OpenStudy (mendicant_bias):

(Lol I don't remember Power Series method at all, so I might be a while on figuring this out)

OpenStudy (mendicant_bias):

Thanks! Looking at some other problems for now (trying to focus on understanding Sturm-Liouville, it's highly unlikely that I'll have to solve a DE using a series method in this upcoming exam), but I will look at that later.

OpenStudy (perl):

sure

OpenStudy (usukidoll):

oh gawd I remember doing this.

OpenStudy (usukidoll):

let me guess.....going crazy on pdes?

OpenStudy (usukidoll):

what they did is that they say suppose lamdba is less than 0 then we have a negative value such that if we have lambda = -k2 and plug it back into the equation, we will see that there are complex roots involved because the equation doesn't factor nicely...sorry for typos. I'm full and bed time soon.

OpenStudy (usukidoll):

in my honest opinion... pdes suck. So glad I passed it, but NEVER AGAIN! Wave equations are easy though.

OpenStudy (mendicant_bias):

I don't see how using the Power Series Method would result in me getting sinh or cosh, still. I understand if exponential functions appeared somehow in the power series solution, that is possible, but that isn't the case, no?

OpenStudy (usukidoll):

hmm I'm trying to remember what I did... I did remember taking three cases when lambda is less than 0, 0, and greater than 0. But, I have never ever used Power Series in them either.